OpenStudy (anonymous):

write the expression in terms of sin and cosine, simply so that no quotients appear in the final expression sin^2x-1/cos(-x)

OpenStudy (campbell_st):

well if you start with sin^2 +cos^2 = 1 then cos^2 = 1 - sin^2 or cos^2 = -1(sin^2 -1) so -cos^2 = sin^2 - 1 make the substitution.... and isn't cos(-x) the same as -cos(x) make that substitution... and it should be easy to simplify now.